MICROWAVE SPECTROSCOPY OF TRANS-ETHYL METHYL ETHER IN THE GROUND STATE

\underline{KAORI KOBAYASHI}, YUSUKE SAKAI, SHOZO TSUNEKAWA, Department of Physics, University of Toyama, 3190 Gofuku, Toyama, 930-8555 Japan; TAIHEI MIYAMOTO, MASAHARU FU-JITAKE and NOBUKIMI OHASHI, Kanazawa University, Japan.

The trans-ethyl methyl ether molecule (CH_{3}CH_{2}OCH_{3}) has two inequivalent methyl group internal rotors which corresponds to the two vibrational motions, ν_{28} and ν_{29}. Due to these internal rotations, a rotational transition could be split into maximum five components. The skeletal torsion (ν_{30}) is another low-lying state (ν_{30}) that interacts with the ν_{28} and ν_{29} modes. The microwave spectra of the trans-ethyl methyl ether molecule in the ν_{28} = 1 ^{a}, ν_{29} = 1^{b} , and ν_{30} = 1^{c} , 2 and 3^{d} have been extensively studied by using Hougen’s tunneling matrix formalism. The microwave spectroscopy in the ground state was studied by several groups ^{e}^{,} ^{f}^{,} ^{g}^{,} ^{h}. The splitting due to the ν_{28} mode (C-CH_{3} internal rotation) is small in the ground state and was not fully resolved in most of the previous studied rotational transitions. In this paper, we report the results of the pulsed nozzle-jet Fourier transform microwave spectroscopy so as to measure the fully resolved spectra. The submillmeter wave spectroscopy was also carried out. Our analysis including the previously reported transitions would be useful for astronomical observations.

aK. Kobayashi, T. Matsui, N. Mori, S. Tsunekawa, and N. Ohashi J. Mol. Spectrosc. \underline{269}, 242 2011.

bK. Kobayashi, T. Matsui, S. Tsunekawa, and N. Ohashi J. Mol. Spectrosc. \underline{255}, 164 2009.

cK. Kobayashi, T. Matsui, N. Mori, S. Tsunekawa, and N. Ohashi J. Mol. Spectrosc.\underline{251}, 301 2008.

dK. Kobayashi, K. Murata, S. Tsunekawa, and N. Ohashi Int. Symposium on Mol. Spectrosc., 65th Meeting TH15 2010.

eM. Hayashi, and K. Kuwada J. Mol. Structure \underline{28}, 147 1975.

fM. Hayashi, and M. Adachi J. Mol. Structure \underline{78}, 53 1982.

gS. Tsunekawa, Y. Kinai, Y. Kondo, H. Odashima, and K. Takagi Molecules \underline{8}, 103 2003.

hU. Fuchs, G. Winnewisser, P. Groner, F. C. De Lucia, and E. Herbst Astrophys. J. Suppl. \underline{144}, 277 2003.